Solve the system of equations
-6x − y = 1
7x + y = 0
x=
y=

Answer by stanbon(75887)

Solve the system of equations
-6x − y = 1
7x + y = 0
------
Add and solve for "x":
x = 1
---
solve for "y":
-6x - y = 1
-6 - y = 1
y = -7
================
Cheers,
Stan H.
========================